OpenStudy (anonymous):

A sequence is defined by the recursive formula f (n + 1) = f(n) – 2. If f(1) = 18, what is f(5)?

OpenStudy (anonymous):

starting with f(1)=18, you get... f(2)=f(1)-2=18-2=16 you can go on till you get f(5) ...

OpenStudy (anonymous):

its 10

OpenStudy (anonymous):

yeah! good work! :)